OpenStudy (anonymous):

If you reveal the top two cards of a deck in succession, what are the chances they both have the same face value (both aces, both kings, etc)? I have a million more like these :(

OpenStudy (anonymous):

would it be 13 choose 1 times something over 52 choose 2?

OpenStudy (anonymous):

so, if the first card can be any number the probability shouldnt matter for the first card

OpenStudy (anonymous):

after the first card is drawn you're left with 51 cards and 3 chances of picking the same card

OpenStudy (anonymous):

so im guessing the answer is 3/51

OpenStudy (anonymous):

do you get what im saying?

OpenStudy (anonymous):

yeah, thank you!

OpenStudy (anonymous):

k

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

there are 52 cards in a deck and you want to pick 2 cards with the same face value so lets say i draw 2 cards and i get 2 aces, then thats one possible combo. There are 4 aces and i want to pick 2 of those 4 aces to get what im looking for which can be represented as \[\left(\begin{matrix}4 \\ 2\end{matrix}\right)\] notice its not limited to aces but it works for 2, 3, 4 .. king so you need to multiply that by 13 \[13\left(\begin{matrix}4 \\ 2\end{matrix}\right)\] remember the binomial theorem which you divide by the total attempts \[\left(\begin{matrix}52 \\ 2\end{matrix}\right)\] \[\frac{ 13\left(\begin{matrix}4 \\ 2\end{matrix}\right) }{ \left(\begin{matrix}52 \\ 2\end{matrix}\right)}\] so your answer should be

OpenStudy (anonymous):

yay! i was almost right, thank you :) 0.0588?

OpenStudy (anonymous):

yea thats what i got